OpenStudy (milrawr):

The density of mercury is 13.6g/mL. How many quarts are in 100 lbs of mercury?

OpenStudy (kevin):

1 lbs = 0.45 kg = 450 g rho = m/V 13.6 = 450/V V = 33.09 mL

OpenStudy (sapphiremoon):

Kevin is right, but don't forget that's the volume for 1 lb of mercury. You'll have to multiply to get it to the volume of 100lb, then convert to find the quarts.